Jose has $20 more than Sara. Their combined money totals $90. Write and solve an algebraic model to determine the amount of money each person has.
Answer by skartikey(21)

let Jose has $J and Sara has $S.
we are given J=S+20 or J-S=20 and J+S=90
by using both equations, we get
2J=110
J=$55
therefore, S=$55-$20=$35